OpenStudy (beccab003):

Solve for x: 2x -3 >(or equal to) 1/2(4x+4) x = all real numbers x has no solution x = 1 I know that it is either the first or second. I think it is the first but I don't know how to know for sure.

OpenStudy (anonymous):

u wanna the steps?

OpenStudy (anonymous):

is it ike this? \(\large 2x -3 \ge\frac{1} {2(4x+4)}\)

OpenStudy (beccab003):

no not quite

OpenStudy (anonymous):

start with expanding \(\large 2x -3 \ge\frac{1} { (8x+8)}\)

OpenStudy (anonymous):

oohh ok :D

OpenStudy (beccab003):

\[2x-3\ge \frac{ 1 }{ 2} (4x+4)\]

OpenStudy (anonymous):

i see :)

OpenStudy (anonymous):

first \(2x-3\ge \frac{ 1 }{ 2} (4x+4)\) \(2x-3\ge ( \frac{ 1 }{ 2} 4x+ \frac{ 1 }{ 2} 4)\) \(2x-3\ge 2x+ 2\)

OpenStudy (anonymous):

so \(2x-3\ge 2x+ 2\) leavs u with \( -3\ge 2\) which is not true for all real :)

OpenStudy (beccab003):

So, it there is no solution for x? That last step you did doesn't make sense because when you move the 2 or the -3 over to one side that you get 5 or -1? you can't have \[-3\]

OpenStudy (beccab003):

\[-3\ge2\]

OpenStudy (beccab003):

that doesn't make sense...

OpenStudy (anonymous):

\(2x-3\ge 2x+ 2\) dear i removed \(\Huge 2x\) \(2x-2x-3\ge 2x-2x+ 2\) \(0-3\ge 0+ 2\) \( -3\ge + 2\)

OpenStudy (anonymous):

see if it make sense to u nw :)

OpenStudy (beccab003):

That makes sense now but you still need to move the 2 over to the other side of the equation to make -1.

OpenStudy (beccab003):

oh I mean -5

OpenStudy (anonymous):

well u got a point its ok if it still as what it is cuz u know \( -3\ge + 2\) is wrong always .. or u cant make it \(-5\ge 0 \) or \(0\ge 5\) any one of them would work

OpenStudy (beccab003):

But it still makes x have no solution. I see

OpenStudy (anonymous):

i told u :) three of them u will have FALSE sentence that make NO SOLUTION TO X

OpenStudy (beccab003):

Thanks!

OpenStudy (anonymous):

np :)

OpenStudy (beccab003):

Can you help me with one more similar one?

OpenStudy (anonymous):

sure

OpenStudy (beccab003):

Solve for x: 5x + 7 > 1/2(10x+14) x = all real numbers x = 1 x has no solution I did the equation and everything ends up canceling out so does that mean that there is no solution?

OpenStudy (anonymous):

yes dear :) 0>0 is false cuz 0=0 so no solution

OpenStudy (beccab003):

so if there is ever a situation like that where there isn't any solution it's just 0=0 then it's x has no solution?

OpenStudy (anonymous):

if u have equation with >= and u end up with cancelling 0>=0 then solution all real number but if there is no = then no solution